OpenStudy (anonymous):

Which mixed number is equivalent to 420%? 4 2/5 4 1/2 4 1/5

OpenStudy (anonymous):

c

OpenStudy (anonymous):

41/5

OpenStudy (anonymous):

C. Every whole number is 100%, so that gives us 400%. Then \[\frac{ 1 }{ 5 }*\frac{ 20 }{ 20 }=\frac{ 20 }{ 100 }=20%\]